### v1

@@ Abstract
Long-form generation is crucial for academic writing papers and repo-level code generation. Despite this, current models, including GPT-4o, still exhibit unsatisfactory performance. Existing methods that utilize preference learning with outcome supervision often fail to provide detailed feedback for extended contexts. This shortcoming can lead to content that does not fully satisfy query requirements, resulting in issues like length deviations, and diminished quality. In this paper, we propose enhancing long-form generation by incorporating process supervision. We employ Monte Carlo Tree Search to gather stepwise preference pairs, utilizing a global memory pool to maintain consistency. To address the issue of suboptimal candidate selection, we integrate external critiques to refine and improve the quality of the preference pairs. Finally, we apply step-level DPO using the collected stepwise preference pairs. Experimental results show that our method improves length and quality on long-form generation benchmarks, with almost lossless performance on general benchmarks across various model backbones.
Figure 1: The above is outcome supervision in long-form generation tasks. Below is LongDPO uses process supervision with a global memory to maintain factual consistency, and external critiques to refine low-reward chosen candidates.
@@ 1 Introduction
Recent advancements in large language models (LLMs) (Zhou et al., 2024; Xiao et al., 2024b; a; Ping et al., 2024) have significantly enhanced their ability to process long text sequences leading models like GPT-4o (OpenAI et al., 2024) can handle context of up to 128K. Despite these advancements, there has been less emphasis on the models’ ability to generate high-quality long-form text outputs. The capability to produce long-form content is essential for various real-world applications, including writing academic papers, novels, and scripts in literature, generating legal contracts in law, and producing repository-level code in technology (Bai et al., 2024b; Wang et al., 2024c). However, many LLMs still struggle to generate content exceeding 2,000 words, highlighting the need for further advancements in this area.
@@
Previous research has investigated methods to extend the output window by constructing long-form training data and utilizing preference learning. For example, Suri (Pham et al., 2024) creates various instructions for the same response and introduces instructional ORPO. LongWriter (Bai et al., 2024b) employs an agent-based pipeline that decomposes ultra-long generation tasks into subtasks to build a long-form dataset, followed by supervised fine-tuning and DPO. These approaches primarily rely on outcome supervision (Lightman et al., 2024) during DPO, which provides feedback on the final result, for long-form generation tasks.
Nevertheless, outcome supervision is not well-suited for long-form generation due to several key challenges. Specifically, outcome supervision means that intermediate steps are not adequately guided, potentially leading to a loss of coherence and inconsistency in the final output Zhang et al. (2024b). As a result, this coarse supervision can produce content that does not fully meet query requirements, leading to problems like length deviations or poor quality Zhang et al. (2024b); Bai et al. (2024b). The absence of fine-grained feedback prevents the model from learning to improve particular aspects of the output.
In this paper, we introduce a novel approach called LongDPO for long-form generation using step-level supervision. Our method is divided into two main components: constructing preference data with stepwise supervision signals and implementing stepwise DPO. Specifically, we employ Monte Carlo Tree Search (MCTS) Browne et al. (2012) with a large language model (LLM) acting as a judge during the evaluation phase to collect stepwise preference pairs. To maintain consistency in the generated text, we incorporate a global memory pool to store factual content from selected nodes during the MCTS search. However, not all chosen candidates in the collected preference pairs are of high quality. To address this issue, we utilize the judge model from the MCTS evaluation phase to provide critiques, followed by critique-augmented generation to refine the chosen candidates. Compared to external critiques, searching for additional solutions is inefficient and may yield only limited performance improvements (Qi et al., 2024). Additionally, relying solely on self-critique, which depends on the model’s inherent capabilities, can result in unstable performance gains (Qi et al., 2024; Zhang et al., 2024c). Given the high-quality stepwise preference pair data, we propose using a stepwise DPO for training to enhance the learning process. As shown in Figure 1, vanilla DPO applies sample-wise supervision directly. Previous work has shown that this approach may result in a less distinct reward margin, making learning more difficult (Lai et al., 2024). In contrast, LongDPO uses fine-grained learning at each step, which may yield better results.
@@
We evaluate long-form generation capabilities with specific length requirements and the ability to follow complex long-form instructions using LongBench-Write-en and LongGenBench (Bai et al., 2024b; Wu et al., 2024c), as well as several general benchmarks such as TruthfulQA (Lin et al., 2022) to assess general task performance. Our method, built on Llama- and Qwen-based backbones, outperforms their DPO versions in long-form generation tasks while maintaining near-lossless performance on general tasks.
@@
  • We propose LongDPO, which differs from traditional outcome supervision by adopting process supervision. LongDPO enables step-wise, more fine-grained learning for long-form text generation.
  • To implement process supervision, we introduce MCTS to construct step-level preference data. Specifically, we utilize a global memory pool to maintain factual consistency and incorporate external critiques to collect better preference pairs.
  • We validate that our method enhances performance in long-form generation tasks while maintaining near-lossless results for general tasks. Meanwhile, we conduct an in-depth analysis to show that the generated responses align more closely with human preferences.
@@ 2 Related Work
Long Context LLMs Some studies explore training-free methods to extend the input context window, such as modifications to the model architecture (Peng et al., 2024; Xiao et al., 2024c), position encoding (Ding et al., 2024; Xiao et al., 2024b), and others. Another approach focuses on training-based techniques (Bai et al., 2024a; Munkhdalai et al., 2024; Fu et al., 2024). Many LLMs can support input context windows of 32K or even over 128K words. However, far fewer are capable of generating outputs exceeding 2K words in length. Recent studies (Pham et al., 2024; Bai et al., 2024b) have employed outcome supervision to extend the output window. Most recently,  Zhang et al. (2024b) proposed LongReward, which is orthogonal to our work. However, in addition to the instruction and response, it requires an additional reference long document as input, which limits its applicability in both outcome and process supervision.
Figure 2: The pipeline of LongDPO. LongDPO incorporates process supervision and MCTS to collect stepwise preference data, where the preference data share the same prefix and only one pair is collected at each layer. During the selection phase, LongDPO uses the global memory pool to filter out candidates that may result in inconsistency, then selects the highest-scoring one as the chosen candidate, with another randomly selected as the rejected candidate. During tree expansion, some chosen candidates may have low rewards, LongDPO uses external knowledge to provide critiques for refinement. Then the collected preference pairs are used for step-level DPO training.
Process Supervision in Preference Learning Recently, scaling inference-time compute has become increasingly popular (Chen et al., 2024b; Setlur et al., 2024; Snell et al., 2024). Process supervision with MCTS can further enhance models’ reasoning abilities (Tian et al., 2024; Zhang et al., 2024d; a). Recent studies (Wang et al., 2024b; Xu et al., 2024) use MCTS in both math and code tasks. In addition to MCTS,  Zhao et al. (2024) also incorporate self-reflection.  Cheng et al. (2024) employ tree search and train a refiner for iterative optimization. In this work, we primarily focus on exploring the potential of process supervision with MCTS in long-form generation.
@@ 3.1.1 Overview
    $UCB_{i}=\alpha\times\sqrt{2\times\ln\left(\frac{N_{i}}{n_{i}}\right)}+v_{i},$ (1)
@@
  • Evaluation: In terms of evaluating each node, we assess each node using the value provided by a reward model, as previous work has demonstrated its effectiveness (Wang et al., 2024b; a). We consider seven principles to evaluate each node. Each principle is rated between 1 and 5, as detailed in Appendix A.1.
@@ 3.1.2 Preference Pair Extraction
To maintain factual consistency as much as possible, we use a global memory pool $M$ to store relevant factual context. After selecting each node $s$, the memory pool $M$ is also updated accordingly. To be specific, after selecting the node $s_{t}$, we extract the factual content of $s_{t}$ using the model $\Theta$. The extracted factual content is denoted as $\{m_{1},m_{2},\dots,m_{k^{\prime}}\}$. After extracting, we first employ $\Theta$ to check the extracted factual contents to ensure they are factually correct as much as possible. We retain only the factual content $\{m_{1},m_{2},\dots,m_{k}\}$ that does not conflict with the internal knowledge of $\Theta$, using the template provided in Appendix A.3. Then, we update the memory correspondingly $M_{t}=M_{t-1}\cup\{m_{1},m_{2},\dots,m_{k}\}$. Next, we will use $M_{t}$ to select the node $s_{t+1}$.
Specifically, after the expansion phase, we visit the nodes in descending order of their average rewards derived using Equation 3. We break the currently visited node $s_{cur}$ into chunks of 128 words, resulting in a sequence of chunks $\{chunk_{1},chunk_{2},\dots,chunk_{j}\}$ and calculate the similarity score using each $m_{k}$ in $M_{t}$ as a query.
$sim_{kj}=E(m_{k})\times E(chunk_{j})^{T},$ (4)
where $sim_{kj}$ is the similarity score, $E(x)$ represents get the embedding of $x$, we use gte-Qwen2-1.5B-instruct [^1] https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct as embedding model. Then, we use the similarity score to filter irrelevant factual content for each $m_{k}$.
$A_{k}=\{chunk_{j}\mid sim_{kj}\geq\delta\},$ (5)
where $\delta$ the similarity threshold is set to 0.8. Finally, we use each $m_{k}$ and its corresponding supported factual content $A_{k}$ to check for any inconsistencies using model $\Theta$ using templates in Appendix A.3. Finally, if no inconsistencies are found, we select $s_{cur}$ as $s_{t+1}$. Otherwise, we will visit the next candidate node without expanding the current one further. For each layer of the tree, we select one pair for preference learning: the node with the highest average reward and no consistency errors is selected as the chosen candidate $s_{win}$, while another node is randomly selected as the rejected candidate $s_{lose}$.
Different from Zhang et al. (2024b), we focus on global information by using the previously selected node as a query, rather than solely focusing on the current local text. On the other hand, we use a more powerful model for a double-check to ensure that the extracted facts are not factually incorrect as much as possible.
@@ 3.2 Chosen Candidates Refinement using Critiques
Collect Data for Critiques Generation $S_{\text{refinement}}$ contains the chosen candidates that need to be refined. Next, we prepare the data for the generation of critiques. Specifically, each data is a triplet $(\text{principle}_{u},s_{sib},s_{win})$, where $\text{principle}_{u}$ is used in the evaluation phase in MCTS to assess the reward of each node, $s_{win}$ is the chosen candidate to be refined, and $s_{sib}$ is the sibling node of $s_{win}$, which serves as an example of refinement as illustrated in Figure 2. Detailed principles are given in Appendix A.1.
@@
Generate critiques Next, we use the reward model $\Theta$ to generate critiques for each triplet using template in Appendix A.2. Figure 4 has shown the main body of the critiques. “Analysis," “Justification," and “Relevant Text" are used to enhance the accuracy of the analysis, while the “Confidence Score" helps assess the model’s confidence in the accuracy of its analysis. “Writing Suggestion" provides recommendations for improvement.
Critique-augmented Generation For each $s_{win}$, we utilize its corresponding critiques $\{z_{1},z_{2},\dots,z_{\lambda}\}$, sorted in descending order by “Confidence Score," to perform critique-augmented generation. Specifically, if $s_{win}$ is selected in layer $t+1$, we rewrite Equation 2 as follows:
$s_{win\_new}=\pi_{\theta}(q\oplus s_{1}\oplus s_{2}\oplus\dots\oplus s_{t}\oplus z_{1}\oplus\dots\oplus z_{\lambda}\ ),$ (7)
where we use each “Writing Suggestion" from $z_{\lambda}$, with a maximum of three. Then, we use the refined data for DPO training.
@@ 3.3 LongDPO Training Objective
In LongDPO, the response $y$ is decomposed into $y=s_{1}\oplus s_{2}\oplus\dots\oplus s_{t}$, where $s_{i}$ represents the $i$-th intermediate result. LongDPO conducts learning at each step. Specifically, for the $(i+1)$-th step, $s_{win}$ is the chosen step, $s_{lose}$ is the rejected step, and $s_{1\sim i}=s_{1}\oplus\dots\oplus s_{i}$ has already been learned. LongDPO aims to maximize the probability of $s_{win}$ and minimize the probability of $s_{lose}$.
$\mathcal{L}_{LongDPO}=-\mathbb{E}_{(q\oplus s_{1\sim i},s_{win},s_{lose})\sim D}\Big{[}\log\sigma\big{(}\\ \beta\log\frac{\pi_{\theta}(s_{win}|q\oplus s_{1\sim i})}{\pi_{ref}(s_{win}|q\oplus s_{1\sim i})}-\beta\log\frac{\pi_{\theta}(s_{lose}|q\oplus s_{1\sim i})}{\pi_{ref}(s_{lose}|q\oplus s_{1\sim i})}\big{)}\Big{]}.$ (9)
Table 1: Evaluation results on LongBench-Write-en. LongWriter-Llama and LongWriter-Qwen represent LongWriter-llama-8B and LongWriter-Qwen2.5-7B, respectively. Our method improves performance across four output length ranges, particularly for outputs exceeding 4K. We have set a random seed to ensure reproducibility.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have markedly improved their capacity to handle long text inputs; however, current models, including GPT-4o, still exhibit unsatisfactory performance in long-form generation. Generating high-quality long-form content still remains a significant challenge. In this paper, we present LongDPO, a novel approach designed to enhance long-form text generation through step-level supervision. By leveraging Monte Carlo Tree Search (MCTS) to collect stepwise preference pairs and employing a global memory pool to maintain factual accuracy, LongDPO effectively mitigates issues such as inconsistencies that are prevalent in long-context LLMs. Furthermore, we integrate critique-augmented generation to refine the selected preference pairs. Following the collection of stepwise preference pairs, we apply stepwise preference learning for fine-grained optimization. Experimental results demonstrate that our method enhances performance on long-form generation benchmarks (e.g. LongBench-Write) while maintaining nearly lossless performance on several general benchmarks. 111  Code and models will be publicly available at https://github.com/pingbowen23/LongDPO.
LongDPO: Unlock Better Long-form Generation Abilities for LLMs via Critique-augmented Stepwise Information
Bowen Ping1, Jiali Zeng2, Fandong Meng2, Shuo Wang3, Jie Zhou2, Shanghang Zhang1 1State Key Laboratory of Multimedia Information Processing, School of Computer Science, Peking University, 2Pattern Recongnition Center, WechatAI, Tencent Inc, 3Dept. of Comp. Sci. & Tech., Tsinghua University, Beijing, China Correspondence:Shanghang Zhang shanghang@pku.edu.cn
Figure 1: The above refers to outcome supervision, which directly provides feedback for extended sequences in long-form generation tasks. Below is LongDPO uses process supervision with a global memory to maintain factual consistency, and external critiques to refine low-reward chosen candidates.
@@ 1 Introduction
Recent advancements in large language models (LLMs) (Zhou et al., 2024; Xiao et al., 2024b, a; Wang et al., 2024b; Ping et al., 2024), have significantly enhanced their capacity to process long text sequences with models like GPT-4o now capable of handling contexts up to 128K tokens (OpenAI et al., 2024; Yang et al., 2025). Despite these strides, there has been less emphasis on the models’ ability to generate better long-form text outputs. The capability to produce long-form content is essential for various real-world applications, including writing academic papers, novels, and scripts in literature, generating legal contracts in law, and producing repository-level code in technology (Bai et al., 2024b; Wang et al., 2024e). However, many LLMs still struggle to generate content exceeding 2,000 words (Pham et al., 2024; Bai et al., 2024b), highlighting the need for further advancements in this area.
@@
Previous research has explored methods to extend the output window by creating long-form training data and leveraging preference learning. For example, Suri (Pham et al., 2024) creates various instructions for the same response and performs outcome-level preference optimization. LongWriter (Bai et al., 2024b) employs an agent-based pipeline that decomposes ultra-long generation tasks into subtasks to build a long-form dataset, followed by supervised fine-tuning and DPO. These approaches primarily rely on outcome supervision (Lightman et al., 2024) during DPO, which provides feedback on the final result, for long-form generation tasks.
Nevertheless, long-context LLMs are more prone to produce responses with issues such as logical inconsistencies, fabricated content, and failure to fully meet query requirements (Zhang et al., 2024b). These challenges make outcome supervision, which directly provides feedback for a long sequence, particularly problematic. In contrast, process supervision involves supervising each intermediate step, which offers more granular and precise feedback. Furthermore, process supervision specifies the exact location of low-quality steps, thereby facilitating the refinement of these steps (Lightman et al., 2024). Consequently, breaking down a long sequence into intermediate steps and supervising these shorter steps could be a more effective strategy.
In this paper, we introduce LongDPO, which enhances long-form generation capabilities through step-level supervision. LongDPO first constructs preference data with stepwise supervision and then performs stepwise learning. Specifically, we use Monte Carlo Tree Search (MCTS) Browne et al. (2012) to collect stepwise preference pairs. Considering that long-context LLMs are prone to generating inconsistent content, leading hallucinations (Zhang et al., 2024b), we incorporate a global memory pool to improve the factual consistency of the selected preference pairs. Additionally, the quality of candidates generated heavily relies on the original model’s inherent capability. Simply searching for candidates is both inefficient and ineffective (Qi et al., 2024). To address this, we propose critique-augmented generation to obtain better candidates for the selected preference pairs.
After gathering the stepwise preference pairs, we propose employing a stepwise DPO for fine-grained learning. As illustrated in Figure 1, traditional DPO applies sample-wise supervision directly, which can lead to a less pronounced reward margin, complicating the learning process (Lai et al., 2024). In contrast, LongDPO utilizes fine-grained learning at each step, which has the potential to produce superior results.
@@
We evaluate long-form generation capabilities using LongBench-Write-en and LongGenBench (Bai et al., 2024b; Wu et al., 2024c), which assess text generation length, quality, and adherence to instructions. Additionally, we use general benchmarks such as TruthfulQA (Lin et al., 2022) to measure overall task performance. Our method, built on Llama- and Qwen-based backbones, outperforms their vanilla DPO versions in long-form generation tasks while maintaining near-lossless performance on general tasks.
@@
  • We introduce LongDPO, which facilitates step-wise, fine-grained learning for long-form text generation.
  • We employ MCTS to create step-level preference data, incorporating a memory pool to enhance factual consistency and external critiques to gather higher-quality preference pairs for long-form generation.
  • The experimental results and in-depth analysis demonstrate the effectiveness of our method in long-form generation tasks.
Figure 2: The pipeline of LongDPO. LongDPO incorporates process supervision and MCTS to collect stepwise preference data. During the selection phase, LongDPO uses the global memory pool to filter out candidates that may result in inconsistency, then selects the highest-scoring one as the chosen candidate, with another randomly selected as the rejected candidate. During tree expansion, LongDPO leverages external critiques only for low-reward chosen candidates. Then the collected preference pairs are used for step-level DPO training.
@@ 2 Related Work
Long Context LLMs Some studies explore to extend the input context window, using training-based methods like (Bai et al., 2024a; Munkhdalai et al., 2024; Fu et al., 2024) and training-free methods, such as (Peng et al., 2024; Xiao et al., 2024c; Ding et al., 2024). Many LLMs can support input context windows of 128K. However, far fewer are capable of generating outputs exceeding 2K words in length. Recent studies (Pham et al., 2024; Bai et al., 2024b) have employed outcome supervision to extend the output window. Most recently,  Zhang et al. (2024b) proposed LongReward, which is orthogonal to our work. However, in addition to the instruction and response, it requires an additional reference long document as input, which limits its applicability in both outcome and process supervision. Another line of exploration in long-text generation, such as hierarchical writing and recurrent prompting (Quan et al., 2024; Xi et al., 2025; Wang et al., 2024c), is orthogonal to our method.
Process Supervision in Preference Learning Recently, scaling inference-time compute has become increasingly popular (Chen et al., 2024b; Setlur et al., 2024; Snell et al., 2024). Process supervision with MCTS can further enhance models’ reasoning abilities (Tian et al., 2024; Zhang et al., 2024d, a). Recent studies (Wang et al., 2024d; Xu et al., 2024) use MCTS in both math and code tasks. In addition to MCTS,  Zhao et al. (2024) also incorporate self-reflection.  Cheng et al. (2024) employ tree search and train a refiner for iterative optimization. In this work, we primarily focus on exploring the potential of process supervision with MCTS in long-form generation.
@@ 3.1.1 Overview
    $\mathrm{UCB_{i}}=\alpha\times\sqrt{2\times\ln\left(\frac{N_{i}}{1+n_{i}}\right)}+v_{i},$ (1)
@@
  • Evaluation: In terms of evaluating each node, we assess each node using the value provided by a reward model, as previous work has demonstrated its effectiveness (Wang et al., 2024d, a). We consider seven principles to evaluate each node. Each principle is rated between 1 and 5, as detailed in Appendix A.1.
@@ 3.1.2 Preference Pair Extraction
We use a global memory pool $M$ storing relevant factual context $\{m_{1},m_{2},\dots,m_{k}\}$ to check consistency before selection. Specifically, after the expansion phase, we visit the nodes in descending order of their UCB scores in Equation 1. We break the currently visited node $s_{cur}$ into contexts of 128 words, resulting in $\{s_{cur_{1}},s_{cur_{2}},\dots,s_{cur_{j}}\}$, each $s_{cur_{j}}$ has 128 words, and calculate the similarity score using each $m_{k}$ in $M_{t}$ as a query.
$\mathrm{{sim}_{kj}}=E(m_{k})\times E(s_{cur_{j}})^{T},$ (4)
where $\rm sim_{kj}$ is the similarity score, $E(x)$ represents get the embedding of $x$, we use gte-Qwen2-1.5B-instruct [^2] https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct as embedding model. Then, we use the similarity score to filter irrelevant context for each $m_{k}$.
$A_{k}=\{s_{cur_{j}}\mid\rm sim_{kj}\geq\delta\},$ (5)
where $\delta$ the similarity threshold is set to 0.8. Finally, we use each $m_{k}$ and its corresponding supported context $A_{k}$ to check for any inconsistencies using model $\Theta$ using templates in Appendix A.3. Finally, if no inconsistencies are found, we select $s_{cur}$ for the next expansion phase. Otherwise, we will visit the next candidate node without expanding the current one further.
After finishing each selection phase, the memory pool $M$ is also updated accordingly. To be specific, after selecting the node $s_{t}$, we extract the factual content of $s_{t}$ using the model $\Theta$ and employ $\Theta$ to verify the extracted factual content to ensure that they are factually correct as much as possible using templates in Appendix A.3. We retain only the factual content $\{m_{1},m_{2},\dots,m_{k^{\prime}}\}$ that does not conflict with the internal knowledge of $\Theta$. Then, we update the memory correspondingly $M_{t}=M_{t-1}\cup\{m_{1},m_{2},\dots,m_{k^{\prime}}\}$.
If memory $M$ is empty, we skip the consistency check and proceed directly to the selection phase and update the memory. When we select $s_{t}$, we only use the factual content stored in $M_{t-1}$, which contains the factual content from the first layer up to the $t-1$ layer.
For each layer of the tree, we select one pair for preference learning: the node with the highest average reward and no consistency errors is selected as the chosen candidate $s_{win}$, while another node is randomly selected as the rejected candidate $s_{lose}$.
@@ 3.2 Chosen Candidates Refinement using Critiques
Collect Data for Critiques Generation $S_{R}$ contains the chosen candidates that need to be refined. Next, we prepare the data for the generation of critiques. Specifically, each data is a triplet $(\text{principle}_{u},s_{sib},s_{win})$, where $\text{principle}_{u}$ is used in the evaluation phase in MCTS to assess the reward of each node, $s_{win}$ is the chosen candidate to be refined, and $s_{sib}$ is the sibling node of $s_{win}$, which serves as an example of refinement as illustrated in Figure 2. Detailed principles are given in Appendix A.1.
@@
Generate critiques Next, we use the reward model $\Theta$ to generate critiques for each triplet using template in Appendix A.2. Figure 3 has shown the main body of the critiques. “Analysis,” “Justification,” and “Relevant Text” are used to enhance the accuracy of the analysis, while the “Confidence Score” helps assess the model’s confidence in the accuracy of its analysis. “Writing Suggestion” provides recommendations for improvement.
Critique-augmented Generation For each $s_{win}$, we utilize its corresponding critiques $\{z_{1},z_{2},\dots,z_{\lambda}\}$, sorted in descending order by “Confidence Score,” to perform critique-augmented generation. Specifically, if $s_{win}$ is selected in layer $t+1$, we rewrite Equation 2 as follows:
$s_{win\_new}=\pi_{\theta}(q\oplus s_{1}\oplus\dots\oplus s_{t}\oplus\dots\oplus z_{\lambda}\ ),$ (7)
where we use each “Writing Suggestion” from $z_{\lambda}$, with a maximum of three. Then, we use the refined data for DPO training.
@@ 3.3 LongDPO Training Objective
In LongDPO, the response $y$ is decomposed into $y=s_{1}\oplus s_{2}\oplus\dots\oplus s_{t}$, where $s_{i}$ represents the $i$-th intermediate result. LongDPO conducts learning at each step. Specifically, for the $(i+1)$-th step, $s_{w}$ is the chosen step, $s_{l}$ is the rejected step, and $s_{1\sim i}=s_{1}\oplus\dots\oplus s_{i}$ has already been learned. LongDPO aims to maximize the probability of $s_{w}$ and minimize the probability of $s_{l}$.
$\mathcal{L}_{LongDPO}=-\mathbb{E}_{(q^{\prime},s_{w},s_{l})\sim D}\Big{[}\log\sigma\big{(}\\ \beta\log\frac{\pi_{\theta}(s_{w}|q^{\prime})}{\pi_{ref}(s_{w}|q^{\prime})}-\beta\log\frac{\pi_{\theta}(s_{l}|q^{\prime})}{\pi_{ref}(s_{l}|q^{\prime})}\big{)}\Big{]},$ (9)
where $q^{\prime}$ represents $q\oplus s_{1\sim i}$, which indicates the query concatenated with the corresponding steps learned up to the $(i+1)$-th step.
@@
Table 1: Evaluation results on LongBench-Write-en. LongWriter-Llama and LongWriter-Qwen represent LongWriter-llama-8B and LongWriter-Qwen2.5-7B. We have set a random seed to ensure reproducibility.